OpenStudy (anonymous):

verify that the points (2,1) (4,0) and (5,7) form the vertices of a right triangle.

OpenStudy (saifoo.khan):

draw it.

OpenStudy (anonymous):

i dont know how to

OpenStudy (saifoo.khan):

Ok, wait.

OpenStudy (anonymous):

wait dont i just have to plot the points?

OpenStudy (saifoo.khan):

can u join me @ twiddla?

OpenStudy (anonymous):

or no

OpenStudy (saifoo.khan):

Oh Yes. & chk whether there is a 90* angle or not.

OpenStudy (anonymous):

it looks like a right angle forma at the point (2,1)

OpenStudy (saifoo.khan):

then its True.

OpenStudy (anonymous):

huh? why?

OpenStudy (saifoo.khan):

coz u hav to verify that does these points make a right angled triangle or not.

OpenStudy (anonymous):

so we dont use the distance formula no where in there?

OpenStudy (saifoo.khan):

there is no such thing written in the ques.

OpenStudy (anonymous):

its and example under my notes for the distance formula

OpenStudy (anonymous):

an*

OpenStudy (anonymous):

u use pythagorus :|

OpenStudy (saifoo.khan):

then u hav to show the working, u should use pythagoras.

OpenStudy (saifoo.khan):

Yeah, he is correct.

OpenStudy (anonymous):

yeah the distance formula

OpenStudy (saifoo.khan):

yup.

OpenStudy (anonymous):

but how? theres three points?

OpenStudy (saifoo.khan):

u hav to take 2, one by one.

OpenStudy (saifoo.khan):

(2,1) (4,0) first these two. (4,0) (5,7) then these two. (2,1) (5,7) finally these two/.

OpenStudy (anonymous):

okay thanks i'll calculuate it and then after i calculate what will indicate i have a vertices?

OpenStudy (saifoo.khan):

then use those in Pythagoras Theorm.

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

huh?\[\sqrt{5},5\sqrt{2},3\sqrt{5}\] to say a^2+b^2=c^2

OpenStudy (anonymous):

those are the answers i got

OpenStudy (saifoo.khan):

Yeah, the formula is correct.

OpenStudy (saifoo.khan):

but what are ABC? name the lenghts.

OpenStudy (anonymous):

oh and sorry i just saw the thing about twiddla i dont have a twiddla account whats twiddla? anything like twitter?

OpenStudy (saifoo.khan):

well there if any1 wants any steps like thing, there v can draw & show it to them.

OpenStudy (saifoo.khan):

http://www.twiddla.com/568385

OpenStudy (anonymous):

may i start another question my computer is freezing

OpenStudy (saifoo.khan):

lool, Ok sure